OpenStudy (anonymous):

Jock decided to swim 10 km up Bega river without leaving the river. He swam 2 km every hour then floated on his back for 10 minutes. Each time he floated, the current dragged him back 1 km. How long did it take Jock to complete the swim?

sam (.sam.):

Do you have the answer?

OpenStudy (anonymous):

it took him 3hours and 20 minutes

sam (.sam.):

d=10km v=2km/h = 1/30 km/min time required=300 min, which is 5hours 5hours =50minutes backwards, 5km backwards=additional 150mins 300+150=450mins or 7.5hours

sam (.sam.):

You check the answer.

sam (.sam.):

Because from the additional 150mins , I'm not sure whether he will move backwards again.

OpenStudy (anonymous):

thanx